SPATIAL DISCRETIZATION OF PARTIAL DIFFERENTIAL EQUATIONS WITH INTEGRALS

ROBERT I. MCLACHLAN

Abstract. We consider the problem of constructing spatial finite difference approximations on a fixed, arbitrary grid, which have analogues of any number of integrals of the partial differential equation and of some of its symmetries. A basis for the space of of such difference operators is constructed; most cases of interest involve a single such basis element. (The “Arakawa” Jacobian is such an element.) We show how the topology of the grid affects the complexity of the operators.